OpenStudy (anonymous):

lim(v→2)(4v^2-8v)/(v-2) please help solve this

OpenStudy (solomonzelman):

\(\large\color{slate}{\displaystyle\lim_{v \rightarrow ~2}\frac{4v^2-8v}{v-2}}\) \(\large\color{slate}{\displaystyle\lim_{v \rightarrow ~2}\frac{4v(v-2)}{v-2}}\)

OpenStudy (solomonzelman):

cancel the `v-2` and plug in 2 for v.

OpenStudy (solomonzelman):

it will be one of those cases,